OpenStudy (anonymous):

Find a non-zero, two-by-two matrix such that: [-6 -5] x [__ __] = [0, 0] [24 20] [__ __] [0, 0] It's a question that has been asked one year ago by someone else but I can't figure out how they get the matrix : [-5 -5] [6 6] When I try to do it. Here is what I get and where I get stuck : -6a -5c = 0 -6b-5d= 0 24a-20c = 0 24b-20d = 0 At this point, I just have NO IDEA how to solve these equations. I am sure it's something really easy to do but I still need help.

OpenStudy (wolfe8):

If you crossed the the matrices correctly and got those equations, solving them simultaneously gets you 0=0?

OpenStudy (anonymous):

Well that's the problem, I have no idea what to do. I'm pretty sure it's something simple that I am missing. Here is what I do : -6a-5c = 0 so a = -(5/6)c Then I replace the a in -6a-5c=0 with the -(5/6)c. I end up with 0=0.. Sorry for my English.

OpenStudy (wolfe8):

I see your problem. When you solve one equation for an unknown, you cannot replace it back in the same equation. Try into another one.

OpenStudy (anonymous):

I just replaced the a in 24b-20a = 0. I end up with c = (-36/25)b. I still have no idea what I need to do to be honest. Do I have to do this for all the equations and then try to figure out a pattern?

OpenStudy (wolfe8):

Now you have c. Replace c in other equations. Keep going until you get a value for any unknown. Then start substituting again to get other values.

OpenStudy (anonymous):

I'm so sorry but I'm still really confused. I never end with a really value when I do that unless I'm missing something. a = -(5/6)c b = -(5/6)d a = (5/8)c b = (5/8)d d = (8/5)b

OpenStudy (wolfe8):

Oh my. I'm sorry but I can't show you step-by-step to do this now because I'm going to bed. But I think I see a mistake in doing the cross of the matrices. Where did your d go?

OpenStudy (anonymous):

Yes I'm sorry I did a mistake on my first post. Here is an explication I found : "Then working out the top left coefficient of your matrix product will give you a relationship between a and c, so you can eliminate one of them (express a in terms of c, in B), and similarly for the top right. B will now just have two unknowns which can take any value. You don't need to worry about the second row of A as it is a multiple of the first row." I am so confused and lost. I remember doing this kind of equations when I was 16 but it has been so long that I feel stupid for not knowing how to simply solve two equations. Thanks for your time!

OpenStudy (wolfe8):

No problem

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

Ok thank you goformit100. I still didn't get an answer on how to solve so I can't close the question for now sorry.

OpenStudy (anonymous):

I will reformulate my question. How to solve these equations : -6a -5c = 0 -6b-5d= 0 24a-20c = 0 24b-20d = 0

OpenStudy (zzr0ck3r):

in the first ROW you have of matrix 1 you have -6 -5 we need -6x-5y=0 two such numbers are x = 5 and y =-6 so we need to make sure if this works for ROW 2 of matrix 1 row two is 24 20 24**5-6*20 = 0 so 5 -6 is the first COLUMN of matrix 2

OpenStudy (zzr0ck3r):

the same argument holds for row 2